Can you tell me the answer to this problem and explain how you got the answer? My father and I have different answers to this problem. he says its 1.85. Here is the problem: 4/7 of the birthday cake was eaten on your birthday. The next day your dad ate 1/2 of what was left. You get to finish the cake, how much is left? Thanks
Answer by ankor@dixie-net.com(12692)

4/7 of the birthday cake was eaten on your birthday. The next day your dad ate 1/2 of what was left. You get to finish the cake, how much is left?
:
7/7 - 4/7 = 3/7 remaining
:
1/2 * 3/7 = 3/14 consumed by Dad
:
14/14 - 8/14 - 3/14 = 3/14 left for you
: